OpenStudy (logan13):

6 inches 12 inches 12 inches 36 inches

OpenStudy (logan13):

https://app159.studyisland.com/images/blank.gif

OpenStudy (logan13):

@AlexandervonHumboldt2

OpenStudy (logan13):

What is the circumference of the circle shown below?

OpenStudy (logan13):

@AlexandervonHumboldt2

OpenStudy (logan13):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

it doesnt load

OpenStudy (alexandervonhumboldt2):

circumfence=ragius*2*pi=diameter*pi

OpenStudy (logan13):

6 in circle

OpenStudy (alexandervonhumboldt2):

cicumfence=6*pi

OpenStudy (logan13):



OpenStudy (alexandervonhumboldt2):

oh so 6 is radius?

OpenStudy (alexandervonhumboldt2):

ok then 6*2*pi=12pi